If 80 cm3 of methane (CH4)
diffuses through a porous
membrane in 20 seconds while 50
cm3 of gas Q diffuse through the
same membrane in 25 seconds,
calculate the molecular mass of
gas Q.
[CH4=16.0]
1
Expert's answer
2020-05-05T03:37:15-0400

R1/R2=√M2/M1. 25*80/50=40

M2=(R1/R2)^2 x M1 = (20/40) x 16 = 4.

He.